OpenStudy (anonymous):

Two panes are headed towards eachother, an are 440 iles apart. If one is going 240 miles an hour and the other 360 mph how far will they be four minutes before meeting?

OpenStudy (anonymous):

To get the time until they meet, you can calculate: x[(240+360)]=400

OpenStudy (anonymous):

Which means they will meet in 48 minutes. To calculate the distance at 4 minutes before they meet, you can see this will be 44 minutes.

OpenStudy (anonymous):

44/60 (because there are 60 minutes in an hour) * (240+360) = 366 2/3

OpenStudy (anonymous):

Ho did you figure it?

OpenStudy (anonymous):

Then you can subtract that from 440: 440-366 2/3 = 33 1/3 miles apart. Look right Anwar? :)

OpenStudy (anonymous):

This is confusng

OpenStudy (anonymous):

I skipped a step to start, so lets go back there. And I mistyped. Sorry. Let me start over. To calculate the time it takes for the planes to meet, we need a period of time that will have the total distance traveled to be 440mi. Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay. To do this, we need an amount of time that, when multiplied by the speed of both planes, equals 440. So we can set it up like this: x*(360+240)=440 (this is where I messed up originally because I said 400) Got that part?

OpenStudy (anonymous):

360+240 is 500

OpenStudy (anonymous):

Right, but do you understand the equation that I posted? x*(360+240)=440, where x is the amount of time it takes for the two planes to meet.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay. Now, we can solve for x: x*(500)=400 Therefore x=440/500 x=22/25

OpenStudy (anonymous):

But we are not done, that is just the time it takes for the two planes to meet. Make sense so far?

OpenStudy (anonymous):

Actually it's 600x=400 :).. But I don't think you need that.

OpenStudy (anonymous):

Oi you are right, thank you. 600x=400 so x=2/3

OpenStudy (anonymous):

Since the speed is in MPH, our x is calculated in hours. If we want 4 minutes before they meet, we should convert the time (x) to minutes by multiplying by 60. We get the time to be 40 minutes. So in 40 minutes the planes will have traveled a total of 440mi and will meet. 4 minutes before that is 36 minutes. At 36 minutes we can calculate the total distance traveled by using the same equation but switching the variables. First, we need to change back to hours by dividing 36 by 60 = .6 Then our x from our earlier equation is .6 and we have: .6*(360+240)=360 So in 36 minutes, which is 4 minutes before the planes meet, the planes have traveled 360 mi towards each other. If they started 440mi apart, after 36 minutes they are 440-360 = 80mi apart.

OpenStudy (anonymous):

Haha, I think there is a slight mistake. the equation should be 600x=440.

OpenStudy (anonymous):

Anyway, I think the problem can be solved in much simpler way, since the planes are not accelerating (a=0). The distance 4 minutes before meeting is the same as the distance they will move in 4 minutes. That's 4(360/60 + 240/60)=4(10)=40 miles.

OpenStudy (anonymous):

Oh. You're so smart. I just like making things complicated.

OpenStudy (anonymous):

Thanks for the compliment.:) You're smart too.